An audio speaker producing a steady sound at an outdoor concert is 20 ft away from you. If you move to a position where the speaker is 63 ft distant, by what factor will the amplitude of the sound change?



1
Expert's answer
2021-08-04T16:29:11-0400

Amplitude is inversely related to distance from the source of generation


Thus


Amplitude is proportional to "\frac{1}{x}"


At first instance,


Amplitude is proportional to "\frac{1}{20}"


At other instance


Amplitude is proportional to "\frac{1}{63}"


Thus


"\frac{initial amplitude}{final amplitude}" = "\frac{63}{20}"